OpenStudy (gabylovesyou):

sqrt(96) + sqrt(150) - sqrt(54)

OpenStudy (gabylovesyou):

@Mertsj

OpenStudy (gabylovesyou):

@Diogo

OpenStudy (gabylovesyou):

is it 6 sqrt(6) ???

OpenStudy (gabylovesyou):

@jhonyy9

OpenStudy (goformit100):

Yes it is 6 sqrt(6) :) Good

jhonyy9 (jhonyy9):

ok but how can you get this result ?

jhonyy9 (jhonyy9):

sqrt(96) + sqrt(150) - sqrt(54) sqrt96 = sqrt(6*16) = 4sqrt6 sqrt150 = sqrt(6*25) = 5sqrt6 sqrt54 = sqrt(9*6) = 3sqrt6 so than you can rewriting your exercise like 4sqrt6 +5sqrt6 -3sqrt6 = sqrt6 +5sqrt6 = 6sqrt6 so than this answer is right sure